From nch@roe.ac.uk Fri Sep 28 11:25:07 2007 Date: Fri, 28 Sep 2007 11:14:36 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Johann Bryant , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Mark Holliman , Peredur Williams , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 28/09/07 Minutes of WFAU VDFS Science Archive meeting: 28th September 2007 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, RSC, NJC, JB, LGR, MAR, JMS Apologies: BCW, JPE, AL, RGM, PMW, ETWS, MSH NB: DONM: 10am, Friday 5th October 2007 in the Plate Library Actions discharged this week: ----------------------------- ACTION: NCH to pester STH/MJI about any revised timescale for receipt of zeropoint recalibration info from CASU. Discharged by MJI, who was in regular contact over the past week and kindly supplied the relevant files on Wednesday - ta. Actions partly discharged but continuing: ----------------------------------------- The following from last time partly done but continue: o) Add in a default row for every detector appearing in every detection table (for schema consistency when querying merged sources and individual detections) - ACTION: RSC & NCH Continues; will get sorted (honestly!) as part of the general ingest DB schema revamp to facilitate rapid generalised photometric/astrometric recalibration after DR3. ACTION: RGM to look through requirements docs to see if any users have mentioned GALEX. Continues; RGM supplied GALEX details; will now check WSA SRAD and VISTA URD for any mention of this. Actions carried forward from 21/09/07 meeting: ---------------------------------------------- ACTION: JB to restore BestDR1 to thutmose from LTO tape at a convenient time (but no rush!). Specific points and new actions: -------------------------------- Project management: Nothing of note this week. WFCAM & VISTA updates: Brief updates on WFCAM (possibly solution to channel edge noise problem) and VISTA (camera performance) were given by NCH and JMS respectively. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: NCH noted that he had tried (and failed) to get the SegueDR6 database backup files transfered from Chigaco in the week; not sure why the Sector SW failed but ETWS is the expert so can look into it on his return from ADASS. WSA Operations: NCH noted that the WSA database has been rationalised in preparation for DR3 release by source and detection filegroup shrinkage; the DB size is now 1.7TB after 0.6TB of free space was recovered and there should be no disk space problems in building the DR3 DB. MAR noted progress with the scripted QC for DR3. Hardware: The power cuts to the site continue, with the last one occuring at 5am on Thursday morning; normality was resumed at around 9am. If this continues, there will be inevitable knock-on delays to the DR3 production schedule. NCH noted that thutmose has been poorly this week with logical volume G: falling off the SCSI bus regularly; suspect either backplane or RAID card problems. Eclipse have been called in and will be here at midday to take a look. Software: RSC reported further work on UDS quality error-bit flagging, work with NCH on the provencance updater and preparations for a new release from the SVN trunk for the curation software. MAR has tweaked the WSA user interface to facilitate flat file access to calibration observations. Everyone else has been at ADASS this week. Survey Data Release: Revised schedule (weeks remaining) in the light of the past week's progress: Weeks What: 1.0 CU8 (photometric zeropoint recalibration) 2.0 QC1 (much work to be done in parallel with previous tho') 0.0 CU5 (diff images for the GPS - after QC, but in parallel with the following and it's very quick anyway) 1.0 Quality bit flagging 3.0 CU7 (source merging from scratch again, unfortunately) 0.0 CU13/14 for DXS/UDS (in parallel with shallow survey CU7s) 2.0 Final CUs. Note that the degree to which some of QC1 can run in parallel is a little uncertain and the above estimates are generally pessimistic; the worst-case scenario of is DR3 release by the end of November. Everything will be done to expedite and arrive sooner. Brief hiatus over the last few days was used productively in shrinking the WSA and updating the processed file provenance. QC1 is progressing well. Eyeball deprecations for the LAS are promised for early next week (electricity supply permitting); GPS will probably be the last to arrive (hopefully within the next week or so). This weekend we plan to update the ZP calibrations with the latest info as supplied earlier in the week by CASU. ACTION: JB, NJC and RSC to make setting of CU8 (photometric ZP recal) a top priority today so it can run over the weekend. Non-survey Data Release: As noted above, MAR has enabled general flat file access to calibration observations by adding in "UKIRT Faint Standards observations" as an option in the archive listing drop-down menu (title was chosen to say exactly what's in the tin, rather than "Calibration" since as we all know the calibration is done wrt to 2MASS, not the UKIRT FS observations). Astrogrid deployment: The team noted that the power outages and poorly state of thutmose have likely impacted the availability of WFAU DSA database services this week. Miscellaneous: Nothing else this week.